# An EFT for anisotropic anti-ferromagnets: gapped Goldstones, pseudo-Goldstones, and phase transitions

### v1

@@ Abstract
We build and discuss a low energy effective field theory for anisotropic anti-ferromagnets in presence of an external magnetic field. Such an effective theory is simple yet rich, and features a number of phenomena such as the appearance of gapped Goldstones, pseudo-Goldstones and a “spin flop” phase transition, all within the regime of validity of the theory. We also discuss in detail, the quantization procedure of the free theory in the presence of a magnetic field, which is made non-trivial by the presence of a single-time derivative term. This class of materials make a precious test field for exotic phenomena in quantum field theory. Moreover, we explicitly perform the matching of the effective theory to the short distance theory of a specific anti-ferromagnet, namely, nickel oxide. The latter is particularly relevant in light of recent proposals of employing this material towards the hunt for light dark matter. As a byproduct of our study, we also re-evaluate the role played by discrete symmetries in magnetic materials, presenting it in a way that is completely consistent with the proper low energy EFT ideology.
@@ I Introduction
Anti-ferromagnets are magnetic materials which, below a certain critical temperature, exhibit long range order but no net magnetization. Contrary to ferromagnets, they were initially thought to have limited practical applications Néel (1971). Recent years, however, have witnessed the development of a plethora of such applications, ranging from electron transport controlled by the spin degrees of freedom (e.g., Baibich et al., 1988; Binasch et al., 1989), the so-called spintronics, the development of quantum sensors (e.g., Lachance-Quirion et al., 2019, 2017), and, recently, light dark matter detection Esposito and Pavaskar (2023); Marsh et al. (2019). Moreover, magnetic materials, both ferromagnets and anti-ferromagnets, are typically regarded as the textbook example for the phenomenon of spontaneous symmetry breaking and the emergence of Goldstone bosons (e.g., Burgess, 2000). These are indeed nothing but the spin collective excitations of these systems, the so-called magnons.
@@
Yet, in reality, the spectrum of most anti-ferromagnets feature tiny but non-zero gaps. This can be attributed to magneto-crystalline anisotropies, which provide small effects of explicit symmetry breaking. This is due to the presence of certain energetically favorable directions for the spins of the material, which are dictated by the underlying crystal structure of the anti-ferromagnet. These anisotropies, along with the exchange interaction, in turn determine the ground state of the system as well its behavior under the application of an external magnetic field.
@@
In this work we extend the well known effective field theory (EFT) for magnons in anti-ferromagnet (e.g., Burgess, 2000; Chandra et al., 1990; Pavaskar et al., 2022) to include these anisotropic effects. In particular, we will consider the case where the system features two distinct anisotropies, the so-called “easy plane” anti-ferromagnets Rezende et al. (2019). The other class of anisotropic anti-ferromagnets, so-called “easy axis”, feature only one anisotropy, but their spectrum can be obtained from that the previous class, simply setting one the parameters to zero. We will show that this extended EFT, while being eventually rather simple, features a number of interesting phenomena, such as the presence of the so-called gapped Goldstones (e.g., Nicolis and Piazza, 2013; Nicolis et al., 2013; Watanabe et al., 2013; Cuomo et al., 2021), their lifting to pseudo-Goldstones, as well as the presence of the so-called “spin flop” phase transition triggered by a varying external magnetic field, where the spin alignment in the ground state suddenly changes its direction (e.g., Rezende et al., 2019). All this, is computable within the regime of validity of the EFT.
@@
Moreover, in the presence of a magnetic field, as we will see, the theory cannot be diagonalized by a local redefinition. Hence it requires a careful quantization procedure, which we address in this work. Finally, we match our EFT to the short distance theory for a particular anti-ferromagnet, namely nickel oxide (NiO). The latter, thanks to its simple structure and spin-interactions, is often considered as the prototypical room-temperature anti-ferromagnet. Indeed, this material has been used to study a number of different processes, such as inelastic light scattering (e.g., Grimsditch et al., 1998; Milano et al., 2004) and magnetic response to terahertz frequencies Sievers III and Tinkham (1963); Satoh et al. (2010). Moreover, as mentioned, it has recently been realized that such a system offers an optimal target to look for the scattering of dark matter with spin-dependent interactions and masses as low as the keV Esposito and Pavaskar (2023), as well as to look for axion-like dark matter with masses in the meV range Catinari et al. (2024).
@@
The goal of this work is also to present everything in a framework which can bridge the linguistic gap that is often present between the standard condensed matter and high energy physics literature. This is particularly important in light of the role that anti-ferromagnets might play in future searches for light dark matter, a genuinely high energy physics question which, however, requires condensed matter tools to be tackled.
@@ II.1 The simplest EFT: the isotropic case
Let us start by reviewing the construction of the standard EFT for anti-ferromagnets, which has been discussed, for example, in Chandra et al. (1990); Burgess (2000); Pavaskar et al. (2022). We will go through this review rather thoroughly, in order to set the stage to the subsequent study, as well as highlight a few points which can make an easier connection between the high and low energy formalisms used to described anti-ferromagnets.
At distances comparable to its lattice spacing, an anti-ferromagnet is essentially a collection of magnetic moments, which anti-align with respect to the neighboring ones. These magnetic moments typically arise from the fundamental spin of the unpaired electrons localized around each atom of the crystal. The exchange and superexchange interactions between electrons pertaining to different atoms give rise to a coupling between these spins. The simplest example of a short distance Hamiltonian describing such a system is given by the Heisenberg model (e.g., Lovesey, 1984),
@@
where $\bm{i}$ and $\bm{j}$ are positions on the lattice, which is usually seen as “bipartite”, i.e. as composed of two sub-lattices, which we label as $\mathds{A}$ and $\mathds{B}$. $\bm{\mathcal{S}}_{\bm{i}}$ is the spin pertaining to a given lattice site, and $J_{\bm{i}\bm{j}}$ are the above mentioned couplings. The Heisenberg model itself is an effective description of the Hubbard model, describing the interactions of electrons on a lattice. For an anti-ferromagnet, all the spins have the same magnitude, $|\bm{\mathcal{S}}_{\bm{i}}|=\mathcal{S}$.
In an anti-ferromagnet, the couplings $J$ are such that the zero-temperature classical ground state, i.e. the state corresponding to the minimum classical energy, [^1] Contrary to ferromagnets, the exact quantum ground state of an anti-ferromagnet is unknown, and it is only approximated by its classical counterpart. Yet, the spin wave theory built out of this latter ground state gives a very accurate description of experimental results (e.g., Lovesey, 1984). corresponds to a configuration where all the spins in the sub-lattice $\mathds{A}$ are aligned parallel to each other, and those in the sub-lattice $\mathds{B}$ are aligned in the opposite direction. The emergence of this ground state is quantified by an order parameter, the Néel vector, given by $\bm{\mathcal{N}}\equiv\sum_{\bm{i}\in\mathds{A}}\bm{\mathcal{S}}_{\bm{i}}-\sum_{\bm{i}\in\mathds{B}}\bm{\mathcal{S}}_{\bm{i}}$. On the classical ground state, the Néel vector acquires an expectation value parallel to the spins, $\langle\bm{\mathcal{N}}\rangle\neq 0$, which can in principle be aligned in any direction. The spin-wave excitations of the system, called magnons, are the long wavelength fluctuations of the order parameter around its equilibrium value.
@@
Consequently, the theory must be invariant under their combined action, which leaves the order parameter unchanged. [^2] In the literature, instead of the discrete rotation we are using here, one often finds a discussion involving the action of a translation by a single lattice site (e.g., Burgess, 2000), which effectively exchanges the two sub-lattices, $\mathds{A}$ and $\mathds{B}$. Here we prefer to phrase everything in terms of the rotation for two reasons: (1) it is better defined at long wavelengths, where one is insensitive to distances comparable to the lattice spacing, and (2) it allows for a unified description of ferromagnets and anti-ferromagnets, as we explain in Appendix A. As we explain more in detail in Appendix A, from this viewpoint, ferromagnets and anti-ferromagnets share the same symmetry breaking pattern, ${\rm SO}(3)\times\mathcal{T}\to{\rm SO}(2)\times(\mathcal{T}\,\mathcal{R}_{\pi})$. What discriminates them is the behavior of the magnetization, i.e. the spin density,. under either of these discrete symmetries. For a ferromagnet it changes sign under time reversal, while for an anti-ferromagnet it does not. Moreover, for a ferromagnet the background magnetization is non-zero, while for an anti-ferromagnet it vanishes. The EFT we present here is built in order to satisfy these requirements.
@@
$\begin{split}\mathcal{L}_{0}={}&\frac{c_{1}}{2}\left[{(\partial_{t}\hat{\bm{n}})}^{2}-v_{\theta}^{2}{(\nabla_{i}\hat{\bm{n}})}^{2}\right]\,,\end{split}$ (3)
all possible additional terms necessarily contain more derivatives and are therefore suppressed in the low energy/long wavelength limit.
Note that by treating time derivatives and spatial gradients separately we assumed the explicit breaking of boost invariance (whether Galileo or Lorentz is irrelevant). This is ultimately not what happens in Nature, as boosts are always spontaneously broken by the underlying lattice, and the ordinary phonons are the corresponding Goldstones bosons necessary to recover the full symmetry, which they realize non-linearly (e.g., Leutwyler, 1997; Nicolis et al., 2015). Our approximation is valid as long as we are not interested in the phonon dynamics and their interplay with magnons. For an EFT including both, see Pavaskar et al. (2022).
In the presence of an external uniform magnetic field, the microscopic Heisenberg Hamiltonian in Eq. (1) is modified by the Zeeman term,
$H_{B}=H_{0}-\mu\sum_{\bm{i}}\bm{B}\cdot\bm{\mathcal{S}}_{\bm{i}}\,,$ (4)
where $\mu$ is the gyromagnetic ratio of the spins. At the level of the EFT, as explained for example in Brauner (2024), this can be included by promoting the time derivative in the Lagrangian $\mathcal{L}_{0}$ to a “covariant” derivative, $\partial_{t}\hat{\bm{n}}\to\partial_{t}\hat{\bm{n}}+\mu\bm{B}\times\hat{\bm{n}}$, i.e.,
$\mathcal{L}_{B}=\frac{c_{1}}{2}\left[{(\partial_{t}\hat{\bm{n}}+\mu\bm{B}\times\hat{\bm{n}})}^{2}\,-{v_{\theta}^{2}}{(\nabla_{i}\hat{\bm{n}})}^{2}\right]\,.$ (5)
To determine the spectrum of this theory, we first need to find the background value for $\hat{\bm{n}}$. This is done by minimizing the Hamiltonian density, taken in the static and homogeneous limit, which reads,
$\begin{split}\mathcal{H}_{B}\big{|}_{\rm{stat.},\,\rm{homog.}}={}&-\mathcal{L}_{B}\big{|}_{\rm{stat.},\,\rm{homog.}}\\ ={}&-\frac{c_{1}}{2}{(\mu\bm{B}\times\hat{\bm{n}})}^{2}\\ ={}&\frac{c_{1}}{2}\mu^{2}B^{2}\left(\hat{n}_{z}^{2}-1\right)\,,\end{split}$ (6)
where we take the magnetic field to be aligned along the $z$-axis, $\bm{B}=B\>\hat{\bm{z}}$, a configuration we will assume in the rest of the work. The Hamiltonian above is clearly minimized by any configuration such that $\hat{n}_{z}=0$. Therefore, while in absence of magnetic field the background value of the order parameter can be aligned along any direction, now this freedom is restricted to the plane perpendicular to the magnetic field itself. We will then take our background value to be, for example, $\langle\hat{\bm{n}}\rangle=\hat{\bm{y}}$.
Following the standard coset construction ideas Callan Jr et al. (1969); Coleman et al. (1969), the magnon fields can be parametrized as local broken transformations around equilibrium. Specifically, the background spontaneously breaks the rotations generated by $S_{1}$ and $S_{3}$, meaning that we can express the magnon fields as $\hat{\bm{n}}(x)=e^{i(\theta^{1}(x)S_{1}+\theta^{3}(x)S_{3})}\cdot\hat{\bm{y}}$. The corresponding quadratic Lagrangian is then,
$\begin{split}\mathcal{L}_{B}={}&\frac{c_{1}}{2}\bigg{[}{\big{(}\dot{\theta}^{1}\big{)}}^{2}-v_{\theta}^{2}{\big{(}\bm{\nabla}\theta^{1}\big{)}}^{2}-\mu^{2}B^{2}{\big{(}\theta^{1}\big{)}}^{2}\\ &+{\big{(}\dot{\theta}^{3}\big{)}}^{2}-v_{\theta}^{2}{\big{(}\bm{\nabla}\theta^{3}\big{)}}^{2}+\mathcal{O}\big{(}\theta^{4}\big{)}\bigg{]}\,,\end{split}$ (7)
@@
$\omega_{q,+}={}$ $\sqrt{\mu^{2}B^{2}+v_{\theta}^{2}q^{2}}\,.$ (8b)
@@
What happens at finite magnetic field? First of all, the existence of the gapless mode is the hallmark of the fact that the background Hamiltonian in Eq. (6) is still degenerate, but now only on the plane perpendicular to the $z$-axis. The gapless mode is the corresponding Goldstone boson. Secondly, the second mode, has a gap that is universal, meaning that it does not depend on the microscopic details of the system, which here are encoded in the effective coefficient $c_{1}$: it is solely dictated by the magnetic field. This is, instead, the hallmark of the so-called gapped Goldstones Nicolis and Piazza (2013); Nicolis et al. (2013); Watanabe et al. (2013); Cuomo et al. (2021), which arise when a system has a finite density for one of its non-Abelian broken generators. This is indeed the case here for the broken charge $S_{3}$. In fact, the background value of the spin density, obtained from the SO(3) Noether current derived from the Lagrangian (5), is given by,
$\langle\bm{s}\rangle=\big{\langle}c_{1}\left[\partial_{t}\hat{\bm{n}}\times\hat{\bm{n}}+\bm{B}-(\bm{B}\cdot\hat{\bm{n}})\,\hat{\bm{n}}\right]\big{\rangle}=c_{1}B\hat{\bm{z}}\,.$ (9)
As another check of this, one can verify that the Lagrangian (5) can be obtained from the Lagrangian (3), valid in absence of magnetic field, by replacing $\hat{\bm{n}}\to e^{-i\mu BtS_{3}}\cdot\hat{\bm{n}}$, which shows that the background under consideration also breaks time translations, but preserves a combination of them with a spin rotation generated by $S_{3}$. In other words, the symmetry breaking pattern in the presence of an external magnetic field is ${\rm SO}(3)\times H\to\bar{H}=H+\mu BS_{3}$. [^4] To keep the notation light, we omit the discrete symmetries. One recognizes $\bar{H}$ to be precisely the modified Hamiltonian used to defined the ground state of a system at finite density for $S_{3}$. Crucially, gapped Goldstones are still exact Goldstone, meaning that they are needed to realize non-linearly the full SO(3) symmetry of the system.

### v2

# An EFT for anisotropic antiferromagnets: gapped Goldstones, pseudo-Goldstones, and phase transitions

@@ Abstract
We build and discuss a low energy effective field theory for anisotropic antiferromagnets in the presence of an external magnetic field. Such an effective theory is simple yet rich and features a number of phenomena such as the appearance of gapped Goldstones, pseudo-Goldstones, and a first order “spin-flop” phase transition, all within the regime of validity of the theory. We also discuss in detail, the quantization procedure of the free theory in the presence of a magnetic field, which is made non-trivial by the presence of a single-time derivative term. This class of materials makes a precious test field for exotic phenomena in quantum field theory. Moreover, we explicitly perform the matching of the effective theory to the short distance theory of a specific antiferromagnet, namely, nickel oxide. The latter is particularly relevant in light of recent proposals of employing this material towards the hunt for light dark matter. As a byproduct of our study, we also re-evaluate the role played by discrete symmetries in magnetic materials, presenting it in a way that is completely consistent with the proper low energy EFT ideology.
@@ I Introduction
Antiferromagnets are magnetic materials which, below a certain temperature, exhibit long range order but no net magnetization. Contrary to ferromagnets, they were initially thought to have limited practical applications [1]. Recent years, however, have witnessed the development of a plethora of such applications, ranging from electron transport controlled by the spin degrees of freedom [e.g., 2, 3], the so-called spintronics, the development of quantum sensors [e.g., 4, 5], and, recently, light dark matter detection [6, 7]. Moreover, magnetic materials, both ferromagnets and antiferromagnets, are typically regarded as the textbook example for the phenomenon of spontaneous symmetry breaking and the emergence of Goldstone bosons [e.g., 8]. These are indeed nothing but the spin collective excitations of these systems, the so-called magnons.
@@
Yet, in reality, the spectrum of most antiferromagnets feature tiny but non-zero gaps. This can be attributed to magneto-crystalline anisotropies, which provide small effects of explicit symmetry breaking. This is due to the presence of certain energetically favorable directions for the spins of the material, which are dictated by the underlying crystal structure of the antiferromagnet. These anisotropies, along with the exchange interaction, in turn determine the ground state of the system as well its behavior under the application of an external magnetic field.
@@
In this work we extend the well known effective field theory (EFT) for magnons in antiferromagnet [e.g., 8, 9, 10] to include these anisotropic effects. In particular, we will consider the case where the system features two distinct anisotropies, the so-called “easy plane” antiferromagnets [11]. The other class of anisotropic antiferromagnets, so-called “easy axis”, feature only one anisotropy, but their spectrum can be obtained from that the previous class, simply setting one of the parameters to zero. We will show that this extended EFT, while being eventually rather simple, features a number of interesting phenomena, such as the presence of the so-called gapped Goldstones [e.g., 12, 13, 14, 15], their lifting to pseudo-Goldstones, as well as the presence of the so-called “spin flop” phase transition triggered by a varying external magnetic field, where the spin alignment in the ground state suddenly changes its direction [e.g., 11]. All this, is computable within the regime of validity of the EFT.
@@
Moreover, in the presence of a magnetic field, as we will see, the theory cannot be diagonalized by a local redefinition. Hence it requires a careful quantization procedure, which we address in this work. Finally, we match our EFT to the short distance theory for a particular antiferromagnet, namely, nickel oxide (NiO). The latter, thanks to its simple structure and spin interactions, is often considered as the prototypical room-temperature antiferromagnet. Indeed, this material has been used to study a number of different processes, such as inelastic light scattering [e.g., 16, 17] and magnetic response to terahertz frequencies [18, 19]. Moreover, as mentioned, it has recently been realized that such a system offers an optimal target to look for the scattering of dark matter with spin-dependent interactions and masses as low as the keV [6], as well as to look for axion-like dark matter with masses in the meV range [20].
@@
The goal of this work is also to present everything in a framework which can bridge the linguistic gap that is often present between the standard condensed matter and high energy physics literature. This is particularly important in light of the role that antiferromagnets might play in future searches for light dark matter, a genuinely high energy physics question which, however, requires condensed matter tools to be tackled. Moreover, an EFT formulation has the great advantage of allowing to neatly disentangle those phenomena that are universal, i.e. solely due to long-distance physics, from those that are, instead, due to a specific short-distance realization of the system.
@@ II The EFT
We now proceed to the construction of our EFT step-by-step, from the simplest instance to the most involved one. Throughout our manuscript, we assume to work at zero temperature. This is partially motivated by the phenomenological reasons highlighted in Ref. [20], but also due to the fact that the inclusion of thermal effects in a general quantum field theory is typically rather non-trivial and beyond the scope of this work.
@@ II.1 The simplest EFT: the isotropic case
Let us start by reviewing the construction of the standard EFT for antiferromagnets, which has been discussed, for example, in Refs. [9, 8, 10]. We will go through this review rather thoroughly, in order to set the stage to the subsequent study, as well as highlight a few points which can make an easier connection between the high and low energy formalisms used to described antiferromagnets.
At distances comparable to its lattice spacing, an antiferromagnet is essentially a collection of magnetic moments, which antialign with respect to the neighboring ones. These magnetic moments typically arise from the fundamental spin of the unpaired electrons localized around each atom of the crystal. The exchange and superexchange interactions between electrons pertaining to different atoms give rise to a coupling between these spins. The simplest example of a short distance Hamiltonian describing such a system is given by the Heisenberg model [e.g., 21],
@@
where $\bm{i}$ and $\bm{j}$ are positions on the lattice, which is usually seen as “bipartite”, i.e., as composed of two sublattices, which we label as $\mathds{A}$ and $\mathds{B}$. $\bm{\mathcal{S}}_{\bm{i}}$ is the spin pertaining to a given lattice site, and $J_{\bm{i}\bm{j}}$ are the above mentioned couplings. For an antiferromagnet, all the spins have the same magnitude, $|\bm{\mathcal{S}}_{\bm{i}}|=\mathcal{S}$.
In an antiferromagnet, the couplings $J$ are such that the zero-temperature classical ground state, i.e., the state corresponding to the minimum classical energy, [^1] Contrary to ferromagnets, the exact quantum ground state of an antiferromagnet is unknown, and it is only approximated by its classical counterpart. Yet, the spin wave theory built out of this latter ground state gives a very accurate description of experimental results [e.g., 21]. corresponds to a configuration where all the spins in the sublattice $\mathds{A}$ are aligned parallel to each other, and those in the sublattice $\mathds{B}$ are aligned in the opposite direction. The emergence of this ground state is quantified by an order parameter, the Néel vector, given by $\bm{\mathcal{N}}\equiv\sum_{\bm{i}\in\mathds{A}}\bm{\mathcal{S}}_{\bm{i}}-\sum_{\bm{i}\in\mathds{B}}\bm{\mathcal{S}}_{\bm{i}}$. On the classical ground state, the Néel vector acquires an expectation value parallel to the spins, $\langle\bm{\mathcal{N}}\rangle\neq 0$, which can in principle be aligned in any direction. The spin-wave excitations of the system, called magnons, are the long wavelength modulations of the order parameter around its equilibrium value.
@@
$\bm{\mathcal{N}}\xrightarrow[]{\mathcal{T}}-\bm{\mathcal{N}}\,,\qquad\bm{\mathcal{N}}\xrightarrow[]{\mathcal{R}_{\pi}}-\bm{\mathcal{N}}\,.$ (2)
Consequently, the theory must be invariant under their combined action, which leaves the order parameter unchanged. [^2] In the literature, instead of the discrete rotation we are using here, one often finds a discussion involving the action of a translation by a single lattice site [e.g., 8], which effectively exchanges the two sublattices, $\mathds{A}$ and $\mathds{B}$, thus changing the sign of the order parameter. Here we prefer to phrase everything in terms of the rotation for two reasons: (1) it is better defined at long wavelengths, where one is insensitive to distances comparable to the lattice spacing, and (2) we believe that, phrased this way, it allows for a unified description of ferromagnets and antiferromagnets, as we explain in Appendix A. The complete symmetry breaking pattern is then ${\rm SO}(3)\times\mathcal{T}\to{\rm SO}(2)\times(\mathcal{T}\,\mathcal{R}_{\pi})$. The EFT we present here is built in order to implement this pattern.
Now, magnons correspond to local modulations of the order parameter away from equilibrium, represented by a field $\hat{\bm{n}}(x)$, with $\langle\hat{\bm{n}}\rangle\propto\langle\bm{\mathcal{N}}\rangle$ and $\hat{\bm{n}}^{2}=1$. [^3] This can be understood from the fact that the coset space ${\rm SO}(3)/{\rm SO}(2)\simeq{\rm S}^{2}$, which is the unit 2-sphere, and $\hat{\bm{n}}$ denotes the unit vector on ${\rm S}^{2}$. With this at hand, the most general low energy EFT for magnons with at most two derivatives is simply given by [8],
@@
where $c_{1}$ is an effective coefficient that can be determined in terms of parameters of the microscopic Hamiltonian, as done in Ref. [6], while $v_{\theta}$ is the magnon propagation speed. All possible additional terms necessarily contain more derivatives and are therefore suppressed in the low energy/long wavelength limit.
Note that, by treating time derivatives and spatial gradients as completely independent from each other, we assumed the explicit breaking of boost invariance (whether Galileo or Lorentz is irrelevant). This is ultimately not what happens in Nature, as boosts are always spontaneously broken by the underlying lattice, and the ordinary gapless acoustic phonons are the corresponding Goldstones bosons necessary to recover the full symmetry, which they realize non-linearly [e.g., 22, 23]. Our approximation is valid as long as we are not interested in the phonon dynamics and their interplay with magnons. For an EFT including both, see Ref. [10].
In the presence of an external uniform magnetic field, H, the microscopic Heisenberg Hamiltonian in Eq. (1) is modified by the Zeeman term,
$H_{\rm H}=H_{0}-\mu\sum_{\bm{i}}\text{{H}}\cdot\bm{\mathcal{S}}_{\bm{i}}\,$, (4)
where $\mu$ is the gyromagnetic ratio of the spins. At the level of the EFT, as explained for example in Ref. [24], this can be included by promoting the time derivative in the Lagrangian $\mathcal{L}_{0}$ to a “covariant” derivative, $\partial_{t}\hat{\bm{n}}\to\partial_{t}\hat{\bm{n}}+\mu\text{\bf H}\times\hat{\bm{n}}$, i.e.,
$\mathcal{L}_{\rm H}=\frac{c_{1}}{2}\left[{(\partial_{t}\hat{\bm{n}}+\mu\text{\bf H}\times\hat{\bm{n}})}^{2}\,-{v_{\theta}^{2}}{(\nabla_{i}\hat{\bm{n}})}^{2}\right]\,.$ (5)
To determine the spectrum of this theory, we first need to find the background value for $\hat{\bm{n}}$. This is done by minimizing the Hamiltonian density, taken in the static and homogeneous limit, which reads
@@
where we take the magnetic field to be aligned along the $z$-axis, $\text{\bf H}=\text{H}\>\hat{\bm{z}}$, a configuration we will assume in the rest of the work for simplicity. (A more general case is reported in Appendix B.) The Hamiltonian above is clearly minimized by any configuration such that $\hat{n}_{z}=0$. Therefore, while in absence of magnetic field the background value of the order parameter can be aligned along any direction, now this freedom is restricted to the plane perpendicular to the magnetic field itself. We will then take our background value to be, for example, $\langle\hat{\bm{n}}\rangle=\hat{\bm{y}}$.
Following the standard coset construction ideas [25, 26], the magnon fields can be parametrized as local broken transformations around equilibrium. Specifically, the background spontaneously breaks the rotations generated by $S_{1}$ and $S_{3}$, meaning that we can express the magnon fields as $\hat{\bm{n}}(x)=e^{i(\theta^{1}(x)S_{1}+\theta^{3}(x)S_{3})}\hat{\bm{y}}$. The corresponding quadratic Lagrangian is then,
[…]
@@
$\omega_{q,+}={}$ $\sqrt{\mu^{2}{\rm H}^{2}+v_{\theta}^{2}q^{2}}\,.$ (8b)
@@
What happens at finite magnetic field? First of all, the existence of the gapless mode is the hallmark of the fact that the background Hamiltonian in Eq. (6) is still degenerate, but now only on the plane perpendicular to the $z$-axis. The gapless mode is the corresponding Goldstone boson. Secondly, the second mode, has a gap that is universal, meaning that it does not depend on the microscopic details of the system, which here are encoded in the effective coefficient $c_{1}$: it is solely dictated by the magnetic field. This is, instead, the hallmark of the so-called gapped Goldstones [12, 13, 14, 15], which arise when a system has a finite density for one of its non-Abelian broken generators. This is indeed the case here for the broken charge $S_{3}$. In fact, the background value of the spin density, obtained from the SO(3) Noether current derived from the Lagrangian (5), is given by
$\langle\bm{s}\rangle=\big{\langle}c_{1}\left[\partial_{t}\hat{\bm{n}}\times\hat{\bm{n}}+\text{\bf H}-(\text{\bf H}\cdot\hat{\bm{n}})\,\hat{\bm{n}}\right]\big{\rangle}=c_{1}{\rm H}\hat{\bm{z}}\,$. (9)
As another check of this, one can verify that the Lagrangian (5) can be obtained from the Lagrangian (3), valid in absence of magnetic field, by replacing $\hat{\bm{n}}\to e^{-i\mu{\rm H}tS_{3}}\cdot\hat{\bm{n}}$, which shows that the background under consideration also breaks time translations, but preserves a combination of them with a spin rotation generated by $S_{3}$. In other words, the symmetry breaking pattern in the presence of an external magnetic field is ${\rm SO}(3)\times H\to\bar{H}=H+\mu\text{\bf H}\cdot\bm{S}$. [^4] To keep the notation light, we omit the discrete symmetries. One recognizes $\bar{H}$ to be precisely the modified Hamiltonian used to defined the ground state of a system at finite density for the operator $\text{\bf H}\cdot\bm{S}$. Crucially, gapped Goldstones are still exact Goldstone, meaning that they are needed to realize non-linearly the full SO(3) symmetry of the system.